Two angels are complementary. One angle is 10 less than three times the other angle. Find the measure of both angles.
Answer by addingup(3677)

Complementary angles are those that add up to 90 degrees.
Let one angle be 3(x-10), just like the problem says. The other angle is just x.
x + 3(x - 10) = 90 Simplify
x + 3x - 30 = 90 Add 30, both sides. Also add x's on the left
4x = 120 Divide both sides by 4
x = 30, and 3(30-10)= 60